OpenStudy (anonymous):

solve the following system of equations. Enter the X coordinate of the solution. round your answer to the nearest tenth. 5x+2y=22 -2x+6y=3

OpenStudy (anonymous):

try to cancel out on of the varibles. for example, you could mult. the first equation by -3 and then when u add the two equations the y values will cancel out and u can solve for x

OpenStudy (anonymous):

then, after u have the x value, plug it into one of the equations to solve for y

OpenStudy (anonymous):

finally, check your work by plugging both values into an equation and solve to see if it comes out correct

OpenStudy (anonymous):

hstuck what did you get for your answer im still not getting it

OpenStudy (texaschic101):

5x + 2y = 22 --->(-3) -2x + 6y = 3 ------------ -15x - 6y = -66 (result of multiplying by -3) -2x + 6y = 3 ------------add -17x = - 63 x = -63/-17 x = 63/17 x = 3.7

OpenStudy (anonymous):

thank you both for the help :)

OpenStudy (anonymous):